Check the continuity of the function

f: R^2to R at(0,0) , where f is defined by

f(x,y)= e^(2x+y) + tan x


1
Expert's answer
2021-06-18T09:24:12-0400
"\lim\limits_{(x,y)\to(0,0)}f(x,y)=\lim\limits_{(x,y)\to(0,0)}(e^{2x+y}+\tan x)"

"=e^{2(0)+0}+\tan (0)=1"

"f(0,0)=e^{2(0)+0}+\tan (0)=1"

Since


"\lim\limits_{(x,y)\to(0,0)}f(x,y)=1=f(0,0),"

then the function "f(x, y)=e^{2x+y}+\tan x" is continuous at "(0, 0)."